OpenStudy (anonymous):

find x and y values: x-12y=-23, x-y=-1

OpenStudy (studygurl14):

x - 12y = -23 --> x = 12y - 23 x - y = -1 --> x = y - 1 12y - 23 = y - 1 Does that make sense? Can you do the rest?

OpenStudy (michele_laino):

please, try using the substitution method

OpenStudy (anonymous):

from the second equation we have -x+y=1 add this equation and the first equation. You will get the value of y and then putting this value of y in any of the equations, get the value of x. @Blu-Girl

OpenStudy (michele_laino):

the correct substitution, is: \[y-1-12y=-23\]

OpenStudy (anonymous):

I come up with 1,2

OpenStudy (michele_laino):

all right!

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (michele_laino):

x=1, and y=2

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Yeah correct.:)